In chemistry class, 14 liters of a 12% alcohol solution must mixed with a 20% solution to get a 13% solution. How many liters of the 20% solutuon are needed?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

In chemistry class, 14 liters of a 12% alcohol solution must mixed with a 20% solution to get a 13% solution. How many liters of the 20% solutuon are needed?
-----
All L liters of 20%
---
14*12 + L*20 = (14*L)*13
Solve for L

Answer by ikleyn(52781)

.
In chemistry class, 14 liters of a 12% alcohol solution must mixed with a 20% solution to get a 13% solution. How many liters of the 20% solution are needed?
~~~~~~~~~~~~`

Let L be the volume, in liters, of the 20% solution to mix.
---
14*12 + L*20 = (14+L)*13


Solve for L.


There is entire bunch of lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
in this site.

Read them and become an expert in solution the mixture word problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook in the section "Word problems" under the topic "Mixture problems".